American Corporation has two equal shareholders, Mr. Freedom and Brave Inc. In addition to their investments in American stock, both shareholders have made substantial loans to American. During the current year, American paid $180,000 interest each to Mr. Freedom and Brave Inc. Assume that American and Brave have 21 percent tax rates, and Mr. Freedom’s marginal tax rate on ordinary income is 37 percent.

  1. Calculate American’s tax savings from deduction of these interest payments and their after-tax cost.
  2. Calculate Brave’s tax cost and after-tax earnings from its receipt of interest income from American.
  3. Calculate Mr. Freedom’s tax cost and after-tax earnings from his receipt of interest income from American.
  1. Recalculate Brave’s tax cost and after-tax earnings assuming its receipt of interest from American is treated as a constructive dividend.
  2. Recalculate Mr. Freedom’s tax cost and after-tax earnings assuming his receipt of interest from American is treated as a constructive dividend.
Answer #1

a.

Tax savings

$37800

After-tax cost of interest

$142200

b.

Tax cost

$37800

After-tax earnings

$142200

c.

Tax cost

$66600

After-tax earnings

$113400

e.

Tax cost

$13230

After-tax earnings

$166770

f.

Tax cost

$27000

After-tax earnings

$153000

Part A

Tax savings = 180000*21% = 37800

After tax cost of interest = 180000-37800 =$142200

Part B

Tax cost = 180000*21% = 37800

After tax earnings = 180000-37800 =$142200

Part C

Tax cost = 180000*37% = 66600

After tax earnings = 180000-66600 =$113400

Part E

Tax cost = 180000*(1-65%)*21% = 13230

After tax earnings = 180000-13230 =$166770

Part F

Tax cost = 180000*15% = 27000

After tax earnings = 180000-27000 =$153000